OpenStudy (asib1214):

A Solution is made by dissolving 25g KCl to make 250ml of solution. If 20ml of this solution is diluted to 250ml, what is the amount concentration of the final solution?

OpenStudy (abmon98):

Number of moles of KCL:25g/74.5g/mol=0.3355moles Change 250ml into dm^3 we get 0.25 Number of moles=Concentration*Volume C=0.3355/0.25=1.34mol/dm^3 Use dilution Law C1*V1=C2*V2 0.25*1.34=25/1000*C2

OpenStudy (asib1214):

i'm getting 0.000335, can you check if it's correct?